# Xing Huang(a), Eva Kopfer(b), Panpan Ren(c) a)

### v3

@@ 3.2 Hyperboundedness for time-homogenenous decoupled SDEs
We present the following condition.
@@
    (3.13) $|b(z,\gamma)-b(\bar{z},\tilde{\gamma})|+\|\sigma(\gamma)-\sigma(\tilde{\gamma})\|_{HS}\leq K_{M}|z-\bar{z}|+K_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma}).$
@@
Let $\tilde{P}_{t}^{\mu}$ be the associated semigroup to (3.18) and $(\tilde{P}_{t}^{\mu})^{\ast}\gamma$ stand for the distribution of the solution to (3.18) with initial distribution $\gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d})$. To derive the hyperboundedness of $\tilde{P}_{t}^{\mu}$, we first investigate the concentration of the invariant probability measure to (3.18).
@@
(3.19) $\sigma(\gamma)\sigma^{\ast}(\gamma)\geq\delta_{2},\ \ \gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
Then for any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, (3.18) has a unique invariant probability measure denoted by $\Phi(\mu)$. Moreover, there exist constants $N>0$ and $\eta_{0}>0$ such that when $K_{I}<\eta_{0}$, $\Phi$ is a map on $\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$ with
@@
(3.20) $\{\Phi(\mu)\}(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})<C_{0},\ \ \mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
@@
(3.21) $|b(0,\mu)|+\|\sigma(\mu)\|_{HS}\leq K_{I}\|\mu\|_{2}+|b(0,\delta_{0})|+|\sigma(\delta_{0})|,\ \ \mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
@@
Moreover, by (3.19), (3.13) and [34, Lemma 3.2], the Harnack inequality for $\tilde{P}_{t}^{\mu}$ with power $2$ holds, i.e. for any $t>0$, there exists a constant $c_{0}>0$ independent of $\mu$ such that
@@
This implies the uniqueness of the invariant probability measure of (3.18), see for instance [35, Theorem 1.4.1].
@@
This means that $\Phi$ is a map on $\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d})$ with
$\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d}):=\left\{\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}):\|\mu\|_{2}^{2}\leq c_{1}(c_{2}-c_{3}K_{I}^{2})^{-1}\right\}.$
(3) Finally, we study the concentration of $\Phi(\mu)$ for any $\mu\in\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d})$. By (3.2), (3.21) and Itô’s formula for $\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}$, for any $\mu\in\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d})$, we derive
@@
$\leq\varepsilon\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}\left\{-c_{2}(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})\text{\rm{d}}t+c_{1}\text{\rm{d}}t+c_{3}K_{I}^{2}\|\mu\|_{2}^{2}\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},\sigma(\mu)\text{\rm{d}}W_{t}\rangle\right\}$
$+\frac{1}{2}\varepsilon^{2}\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}|\sigma(\mu)^{\ast}[\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu}]|^{2}$
$\leq\varepsilon\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}\left\{(-c_{2}+\varepsilon c_{4})(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})\text{\rm{d}}t+c_{5}\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},\sigma(\mu)\text{\rm{d}}W_{t}\rangle\right\}$
@@
$\varepsilon\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}\left\{(-c_{2}+\varepsilon c_{4})(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})+c_{5}\right\}\leq c_{6}(\varepsilon)-c_{7}(\varepsilon)\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}.$
This together with (3.15) yields $\frac{1}{t}\int_{0}^{t}\{(\tilde{P}_{s}^{\mu})^{\ast}\delta_{0}\}(\text{\rm{e}}^{\varepsilon C_{\psi}^{-1}|\cdot|^{2}})\text{\rm{d}}s\leq\frac{c_{6}(\varepsilon)t+1}{c_{7}(\varepsilon)t}.$ Thus, the assertion (3.20) is derived by following the line to derive (3.24). ∎
The Harnack inequality (3.25) is not as sharp as the one in (2.2). In the following theorem, we will first derive a sharper Harnack inequality, which implies the hyperboundedness of $\tilde{P}_{t}^{\mu}$ for large enough $t$.
@@
(3.26) ×$\quad\times\text{\rm{e}}^{c_{0}C_{1}\frac{C_{\psi}^{2}}{\theta}+c_{0}C_{\psi}^{2}\text{\rm{e}}^{-\frac{\theta}{C_{\psi}}(t-1)}|z-\bar{z}|^{2}},\ \ t\geq 1,f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{2d}),z,\bar{z}\in\mathbb{R}^{2d},\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$
holds for some constant $C_{1}>0$ depending on $r,r_{0},\theta,R,K_{M}$. Consequently, if $K_{I}$ is small enough, then there exist constants $t_{0}>0,\tilde{C}>0$ such that the hyperboundedness of $\tilde{P}_{t_{0}}^{\mu}$ holds, i.e.
@@
Taking expectation in (3.2) and using the semigroup property and Jensen’s inequality, we deduce (3.4). Repeating the same argument to derive (2.26) from (2.2) and (2.20), we gain (3.27) from (3.20) and (3.4). ∎
3.3 Log-Sobolev and Talagrand inequality for $\Phi(\mu)$
@@
    (3.30) $\Phi(\mu)(f^{2}\log f^{2})\leq C_{L}(\Phi(\mu))\Phi(\mu)(|\nabla f|^{2}),\ \ \Phi(\mu)(f^{2})=1,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d})$
@@
  2. 2.
    If moreover, for any $R>0$ and $N>0$, there exist positive constants $C_{1}$ and $C_{2}$ depending on $R$, $N$ such that
@@
    $\Phi(\mu)(f^{2}\log f^{2})\leq\bar{C}\Phi(\mu)(|\nabla f|^{2}),$
    (3.33) $\qquad\quad\Phi(\mu)(f^{2})=1,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}),$
@@
Combining the hyperboundedness (3.27) and the log-Sobolev inequality (3.8) with $s=0$ and $\tilde{P}_{t}^{\mu}$ replacing $P_{s,t}^{\mu}$, we obtain a defective log-Sobolev inequality by Lemma 2.5(1), i.e. there exist constant $\tilde{C}_{1},\tilde{C}_{2}>0$ independent of $\mu$ such that
@@
[…]
The irreducibility of $\mathscr{E}(f,f)=\Phi(\mu)(|\nabla f|^{2})$ implies (3.30) by [37, Corollary 1.3]. This further implies (3.31) according to [27]. Finally, by (3.3), (3.20) and (3.32), the proof is completed by Lemma 2.5(2). ∎
Remark 3.6.
The drift $b$ in Theorem 3.5 is allowed to be partially dissipative, which improves the existing results in [30] where $b$ is assumed to be uniformly dissipative. In contrast to the results in [32] and [4], the drift is allowed to be not of gradient-type. We should also point out that different from the non-degenerate case, the log-Sobolev inequality (3.30) does not imply the exponential convergence of $\tilde{P}_{t}^{\mu}$ in relative entropy since the Dirichlet form is $\Phi(\mu)(|\nabla f|^{2})$ instead of $\Phi(\mu)(|\nabla^{(2)}f|^{2})$.
3.4 Exponential ergodicity of (3.12)
@@ 3.4.1 Uniformly dissipative case
(3.37) in Theorem 3.7 extends the results in [30, Theorem 2.4] and [17, Theorem 4.2], where $b$ is assumed to be of gradient-type, i.e. $b(x,y,\mu)=-y-x-\nabla V(x,\mu)$.
@@
By (C) with $R=0$, it follows from [17, Theorem 4.2] that (3.12) has a unique IPM $\bar{\mu}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ satisfying (3.36). Since $\Phi(\bar{\mu})=\bar{\mu}\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$, (3.31) implies the following Talagrand inequality
@@
By [17, Theorem 3.1], there exists a measurable function $c\colon(0,\infty)\rightarrow(0,\infty)$ such that
@@ 3.4.2 Partially dissipative case
In the distribution independent case, [32, Theorem 39] derives the exponential ergodicity in relative entropy by means of the log-Sobolev inequality for the explicit invariant probability measure on the product space $\mathbb{R}^{2d}$. In Theorem 3.9, we offer an explicit partially dissipative condition under which the log-Sobolev inequality holds. Moreover, the authors in [14] adopt the log-Sobolev inequality and log-Harnack inequality for mean field interacting particle system as well as propagation of chaos to deduce the exponential ergodicity in $L^{2}$-Wasseretein distance and mean field entropy.
@@
$\rho_{\mu}(x,y):=\frac{\text{\rm{d}}\Phi(\mu)}{\text{\rm{d}}x\text{\rm{d}}y}=Z_{0}\exp\left\{-\frac{1}{2}|y|^{2}-V(x)-\int_{\mathbb{R}^{2d}}W(x,z)\mu(\text{\rm{d}}z)\right\}\text{\rm{d}}x\text{\rm{d}}y,$
@@
$Z_{0}=\int_{\mathbb{R}^{2d}}\exp\left\{-\frac{1}{2}|y|^{2}-V(x)-\int_{\mathbb{R}^{2d}}W(x,z)\mu(\text{\rm{d}}z)\right\}\text{\rm{d}}x\text{\rm{d}}y.$
Since $V$ and $W$ are Lipschitz continuous, we conclude that for any $R>0$ and $N>0$, there exist constant $C_{1}$ and $C_{2}$ depending on $R$, $N$ such that
@@
So, by Theorem 3.5, when $K_{I}$ is small enough, we obtain (2), which combined with [32, Theorem 39] implies that we can find constants $c,\lambda>0$ such that
@@
So, when $K_{I}<\min(\eta_{0},\eta_{1})$, we can derive from the Banach fixed theorem that (3.44) has a unique invariant probability measure $\bar{\mu}\in\hat{\mathscr{P}}_{2}(\mathbb{R}^{2d})$.
@@
This together with (3.4.2) for $\mu=\bar{\mu}$ and Gronwall’s inequality yield

### v4

@@ 3.2 Time-homogeneous decoupled SDEs
Let $\tilde{P}_{t}^{\mu}$ be the associated semigroup to (3.13) and $(\tilde{P}_{t}^{\mu})^{\ast}\gamma$ stand for the distribution of the solution to (3.13) with initial distribution $\gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d})$. To derive the hyperboundedness for time-homogenenous decoupled SDE (3.13), we will adopt Wang’s Harnack inequality. Note that in the present degenerate case, the Lipschitz continuity of $b$ in the spatial variable is required to ensure Wang’s Harnack inequality so that we present the following condition.
@@
[…]
@@
3.2.1 Invariant probability measure
In this part, we investigate the existence, uniqueness and concentration of the invariant probability measure to (3.13).
@@ 3.2.1 Invariant probability measure
(3.17) $\sigma(\gamma)\sigma^{\ast}(\gamma)\geq\delta_{2}I_{d\times d},\ \ \gamma\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
  1. $(1)$
    Then for any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$, (3.13) has a unique invariant probability measure denoted by $\Phi(\mu)$.
  2. $(2)$
    Moreover, there exists a constant $\eta_{0}>0$ such that when $K_{I}<\eta_{0}$, we can find a constant $N>0$ such that $\Phi$ is a map on $\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$ with
@@ 3. $(3)$
    (3.18) $(\Phi(\mu))(\text{\rm{e}}^{\varepsilon|\cdot|^{2}})<C_{0},\ \ \mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
@@
(3.19) $|b(0,\mu)|+\|\sigma(\mu)\|_{HS}\leq K_{I}\|\mu\|_{2}+|b(0,\delta_{0})|+\|\sigma(\delta_{0})\|_{HS},\ \ \mu\in\mathscr{P}_{2}(\mathbb{R}^{2d}).$
@@
Moreover, by (3.17), (3.14) and [45, Lemma 3.2], Wang’s Harnack inequality for $\tilde{P}_{t}^{\mu}$ with power $2$ holds, i.e. for any $t>0$, there exists a constant $c_{0}>0$ independent of $\mu$ such that
@@
This implies the uniqueness of the invariant probability measure of (3.13), see for instance [46, Theorem 1.4.1(3)].
@@
Let $N=c_{1}(c_{2}-c_{3}K_{I}^{2})^{-1}$. This means that $\Phi$ is a map on $\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$ with
(3) Finally, we study the concentration of $\Phi(\mu)$ for any $\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$. By (3.2.1), (3.19) and Itô’s formula for $\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}$, for any $\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$, we derive
@@
$\leq\varepsilon\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}\left(-c_{2}(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})\text{\rm{d}}t+c_{1}\text{\rm{d}}t+c_{3}K_{I}^{2}\|\mu\|_{2}^{2}\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},\sigma(\mu)\text{\rm{d}}W_{t}\rangle\right)$
$\quad+\frac{1}{2}\varepsilon^{2}\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}|\sigma(\mu)^{\ast}[\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu}]|^{2}$
$\leq\varepsilon\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}\left((-c_{2}+\varepsilon c_{4})(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})\text{\rm{d}}t+c_{5}\text{\rm{d}}t+\langle\tilde{Y}_{t}^{\mu}+rr_{0}\tilde{X}_{t}^{\mu},\sigma(\mu)\text{\rm{d}}W_{t}\rangle\right)$
@@
$\varepsilon\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}\left((-c_{2}+\varepsilon c_{4})(|\tilde{X}_{t}^{\mu}|^{2}+|\tilde{Y}_{t}^{\mu}|^{2})+c_{5}\right)\leq c_{6}(\varepsilon)-c_{7}(\varepsilon)\text{\rm{e}}^{\varepsilon\tilde{\psi}(\tilde{X}_{t}^{\mu},\tilde{Y}_{t}^{\mu})^{2}}.$
This together with (3.16) yields $\frac{1}{t}\int_{0}^{t}\big(\tilde{P}_{s}^{\mu})^{\ast}\delta_{0}\big)wofangni(\text{\rm{e}}^{\varepsilon C_{\psi}^{-1}|\cdot|^{2}})\text{\rm{d}}s\leq\frac{c_{6}(\varepsilon)t+1}{c_{7}(\varepsilon)t}.$ Thus, (3.18) is derived by following the line to derive (3.22). ∎
3.2.2 Wang’s Harnack inequality and hyperboundedness
The Harnack inequality (3.23) is not as sharp as the one in (2.29). In the following theorem, we will first derive a sharper Wang’s Harnack inequality, which implies the hyperboundedness of $\tilde{P}_{t}^{\mu}$ for large enough $t$. Note that the proof of (3.4) is different from that of (2.29).
@@ 3.2.2 Wang’s Harnack inequality and hyperboundedness
(3.24) $\quad\times\text{\rm{e}}^{c_{0}C_{1}\frac{C_{\psi}^{2}}{\theta}+c_{0}C_{\psi}^{2}\text{\rm{e}}^{-\frac{\theta}{C_{\psi}}(t-1)}|z-\bar{z}|^{2}},\ \ t\geq 1,f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{2d}),z,\bar{z}\in\mathbb{R}^{2d},\mu\in\mathscr{P}_{2}(\mathbb{R}^{2d})$
holds for some constant $C_{1}>0$ depending on $r,r_{0},\theta,R,K_{M}$. Consequently, if $K_{I}$ is small enough, then there exist constants $t_{0}>0,\tilde{C}>0$ such that
@@
Taking expectation in (3.2.2) and using the semigroup property and Jensen’s inequality, we deduce (3.4). Repeating the same argument to derive (2.36) from (2.29) and (2.24), we gain (3.25) from (3.18) and (3.4). ∎
3.2.3 Log-Sobolev inequality
@@ 1. $(1)$
    $\Phi(\mu)(f\log f)-\Phi(\mu)(f)\log\Phi(\mu)(f)$
    (3.28) $\leq C_{L}(\Phi(\mu))\Phi(\mu)(|\nabla f^{\frac{1}{2}}|^{2}),\ \ 0<f\in C_{0}^{\infty}(\mathbb{R}^{2d})$
@@
  2. $(2)$
    If moreover, for any $R>0$ and $N>0$, there exist positive constants $C_{1}$ and $C_{2}$ depending on $R$, NN such that
@@ 2. $(2)$
    $\Phi(\mu)(f\log f)-\Phi(\mu)(f)\log\Phi(\mu)(f)$
    (3.31) $\leq\bar{C}\Phi(\mu)(|\nabla f^{\frac{1}{2}}|^{2}),\ \ 0<f\in C_{0}^{\infty}(\mathbb{R}^{2d}),\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}),$
@@
Remark 3.6.
The drift $b$ in Theorem 3.5 is allowed to be partially dissipative, which improves the existing results in [41] where $b$ is assumed to be uniformly dissipative. In contrast to the results in [43] and [4], the drift is allowed to be not of gradient-type.
@@
(1) Combining the hyperboundedness (3.25) and the log-Sobolev inequality (2.6) with $s=0$ due to Theorem 3.1 and $\tilde{P}_{t}^{\mu}$ replacing $P_{s,t}^{\mu}$, we obtain a defective log-Sobolev inequality by Lemma 2.10(1), i.e. there exist constant $\tilde{C}_{1},\tilde{C}_{2}>0$ independent of $\mu$ such that
@@
(3.33) $\leq\tilde{C}_{1}\Phi(\mu)(|\nabla f^{\frac{1}{2}}|^{2})+\tilde{C}_{2},\ \ 0<f\in C_{0}^{\infty}(\mathbb{R}^{d}),\Phi(\mu)(f)=1,\mu\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d}).$
The irreducibility of $\mathscr{E}(f,f)=\Phi(\mu)(|\nabla f|^{2})$ implies ($(1)$) by [48, Corollary 1.3]. ($(1)$) further implies (3.29) according to [37].
(2) By (3.2.3), (3.18) and (3.30), the proof is completed by Lemma 2.10(2) and [37]. ∎
3.3 Exponential ergodicity of McKean-Vlasov SDEs
@@ 3.3.1 Uniformly dissipative case
(3.35) in Theorem 3.7 extends the results in [41, Theorem 2.4] and [25, Theorem 4.2], where $b$ is assumed to be of gradient-type, i.e. $b(x,y,\mu)=-y-x-\nabla V(x,\mu)$.
@@
By (C) with $R=0$, it follows from [25, Theorem 4.2] that when $K_{I}$ is small enough, (3.10) has a unique IPM $\bar{\mu}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ satisfying (3.34). Since $\Phi(\bar{\mu})=\bar{\mu}\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$, (3.29) implies the following Talagrand inequality
@@
By [25, Theorem 3.1], there exists a measurable function $c\colon(0,\infty)\rightarrow(0,\infty)$ such that
@@ 3.3.2 Partially dissipative case
In the distribution independent case, [43, Theorem 39] derives the exponential ergodicity in relative entropy by means of the log-Sobolev inequality for the explicit invariant probability measure on the product space $\mathbb{R}^{2d}$. In Theorem 3.9, we offer an explicit partially dissipative condition under which the log-Sobolev inequality holds. Moreover, the authors in [22] adopt the uniform in NN(number of particles) log-Sobolev inequality and log-Harnack inequality for mean field interacting particle system as well as propagation of chaos to deduce the exponential ergodicity in $L^{2}$-Wasserstein distance and mean field entropy. Different from [22], in Theorem 3.9, we derive the exponential ergodicity in relative entropy and instead of starting from the mean field interacting particle system as in [22], we begin by investigating the time-homogeneous decoupled SDEs.
@@
$\rho_{\mu}(x,y):=\frac{\text{\rm{d}}\Phi(\mu)}{\text{\rm{d}}x\text{\rm{d}}y}=Z_{0}\exp\left(-\frac{1}{2}|y|^{2}-V(x)-\int_{\mathbb{R}^{2d}}(W(x,z)-W(0,z))\mu(\text{\rm{d}}z)\right),$
@@
$Z_{0}=\int_{\mathbb{R}^{2d}}\exp\left(-\frac{1}{2}|y|^{2}-V(x)-\int_{\mathbb{R}^{2d}}(W(x,z)-W(0,z))\mu(\text{\rm{d}}z)\right)\text{\rm{d}}x\text{\rm{d}}y.$
By (3.38), we can find a constant $C_{W,V}>0$ such that
$|\nabla_{x}W(x,z)|\leq C_{W,V}(1+|x|+|z|),$
$|W(x,z)-W(0,z)|\leq C_{W,V}(1+|x|+|z|)|x|,$
$|V(x)|\leq C_{W,V}(1+|x|^{2}).$
So, we conclude that for any $R>0$, there exist constants $C_{1}$ and $C_{2}$ depending on $R$, NN such that
@@
So, by Theorem 3.5, when $K_{I}$ is small enough, we obtain ($(2)$), which combined with [43, Theorem 39] implies that we can find constants $c,\lambda>0$ such that
@@
So, when $K_{I}<\min(\eta_{0},\eta_{1})$, we can derive from the Banach fixed theorem that (3.42) has a unique invariant probability measure $\bar{\mu}\in\hat{\mathscr{P}}_{2,N}(\mathbb{R}^{2d})$.
@@
This together with (3.3.2) for $\mu=\bar{\mu}$ and Gronwall’s inequality yields